answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Sample space of tossing a coin 6 times?

{hhhhhh, hhhhht, hhhhth, hhhhtt, hhhthh, hhhtht, hhhtth, hhhttt, hhthhh, hhthht, hhthth, hhthtt, hhtthh, hhttht, hhttth, hhtttt, hthhhh, hthhht, hthhth, hthhtt, hththh, hththt, hthtth, hthttt, htthhh, htthht, htthth, htthtt, httthh, htttht, htttth, httttt, thhhhh, thhhht, thhhth, thhhtt, thhthh, thhtht, thhtth, thhttt, ththhh, ththht, ththth, ththtt, thtthh, thttht, thttth, thtttt, tthhhh, tthhht, tthhth, tthhtt, tththh, tththt, tthtth, tthttt, ttthhh, ttthht, ttthth, ttthtt, tttthh, ttttht, ttttth, tttttt}

Edit: It removed all my capitalization, but everything should be capitalized.

A sack contains red and blue marbles The ratio of red marbles to blue marbles is 43 If there are 16 red marbles in the sack how many blue marbles are in the sack?

This problem doesn't lead to a whole number solution. Are you sure you copied it correctly? Maybe you meant the ratio is 4 to 3. In that case:

red/blue = 4/3

blue/red = 3/4

blue = (3 * red) / 4 = (3 * 16) / 4 = 12 blue marbles

9 In how many different ways can you arrange 7 books on a shelf?

Consider one placement at a time. The first book on the shelf could be any of seven. The second could be any of the remaining six. This continues with the rest. The total combinations are:

7! = 7*6*5*4*3*2*1 = 5040

When do you use uniform distribution?

When, over a given range, the probability that a variable in question lies within a particulat interval is equal to the size of that interval as a proportion of the range.

In statistics which is better large sample size or small sample size?

Generally, the larger the sample the more reliable the results.

Example:

If you flipped a coin twice and got heads both times you could say the coined is biased towards heads.

However, if you repeat the experiment 100 times your results will be a lot more reliable.

What do you win if the odds are 5 to 2?

When the odds are 5 to 2, if you bet $2 on something and win then you will receive $7 The $2 you bet and the $5 you won.

What is a probability of drawing a club and a 7?

drawing a club

total club =13

P(Club)= 13/52

drawing a 7

total 7 = 4

P(7) = 4/52

P(a club and a 7) = P(Club) + P(7) as events are mutualy exculisve

P(a club and a 7) = 13/52+4/52

P(a club and a 7) = 17/52

If you roll a fair die what is the theoretical probability of getting a 5?

Since there are 6 numbers on a die (1-6), then the probability of rolling a 5 would be 1 out of 6.

What is the probability of having a girl versus having a boy?

The probability of having a girl versus a boy is 1/2 because there is two things you have a chance of getting and you can only get one or the other.

What is the definititon of a discrete random variable?

The discrete random variable almost always arises in connection with counting.

How do you do a probability?

Number of useful outcomes over number of possible outcomes and simplify it if you can. Imagine you want an even number and you roll a die. There are 6 possible outcomes and three of them are useful outcomes (outcomes we want).

3

6

Simplify it and you get

1

2

You roll a standard number cube and get a multiple of 4?

You got a one or a four. One is the zeroeth multiple of 4. Four is the first multiple of four. Any other result, such as 8, and you are not rolling a standard die.

What is a biased spinner?

ok I'm sure not many people will be asking this but a biased spinner is basically a spinner that is unfair or not accurate

What is standard normal variate?

It's the same as a z-Transformation.

for all xi: (xi-mean(x)) / std(x)

What is the probability of at least one birthday match among a group of 41 people?

The probability of at least 1 match is equivalent to 1 minus the probability of there being no matches.

The first person's birthday can fall on any day without a match, so the probability of no matches in a group of 1 is 365/365 = 1.

The second person's birthday must also fall on a free day, the probability of which is 364/365

The probability of the third person also falling on a free day is 363/365, which we must multiply by the probability of the second person's birthday being free as this must also happen. So for a group of 3 the probability of no clashes is (363*364)/(365*365).

Continuing this way, the probability of no matches in a group of 41 is (365*364*363*...326*325)/36541

This can also be written 365!/(324!*36541)

Which comes to 0.09685...

Therefore the probability of at least one match is 1 - 0.09685 = 0.9032

So the probability of at least one match is roughly 90%

Trending Questions
How many different arrangements can be made with the letters in the word Iowa? How do you take off an spinner from an fan? Uses of probability? What is the full form of ace in a pack of card? What is18 divided by 1? How many different 5 persons committees can be formed from a group of 8 women and 6 men with each committee containing at least 3 women? What percents of cards are diamonds in a standard deck of cards? What is the probability of getting a perfect score in a three-question true or false quiz? What is the possible outcome of tossing 10 coins? What is the probability of getting four heads when flipping four identical coins at once? Calculate the empirical probability of a pair in a 5 card hand of poker? What is the exponential growth formula? What are the percent of chance of rolling a six twice in a row on one six sided dice? What does stringent mean? Taking into account optimal sampling strategy and considering cost implications how do you determine the most appropriate sampling method to use? Lamone is in a group of 10 students on waiting list for a school trip There is space for 2 more students on the tripso Mr.Baraka plansto randomly select 2 names from a hatWhat is the probability? How many favorable outcomes would be possible if the number of total possible outcomes is increased to 24? How many people would you have to have in a group so that there is a probability of at least 0.6 that at least two of them will have the same birthday? What is the probability of drawing a spade out a deck of 52 playing cards? How many cards are diamond or a four in a 52 deck of regular cards?